Acquirer firm plans to launch a takeover of Target firm. The manager of Acquirer indicates that the deal will increase the free cash flow of the combined business by $13.6m per year forever. The beta of the combined firm is 1.2, market portfolio return is 12% and risk free interest rate is 4%. Firms Involved in the Takeover Acquirer Target Assets ($m) 6000 800 Debt ($m) 2000 300 Number of shares (m) 80 20 (i). Calculate the value of synergy of the deal. (6 marks) 10 (ii). Calculate the offer price at which Acquirer shares the total synergy as in (i) equally with Target
Answer #1

Rate of Return = Rf + b (Rm - Rf ) =

Given Beta = 1.2,Market Return = 12%, Risk Free Interest = 4%

Rate of Return = 4 + 1.2 (12 - 4 ) = 13.6%

Cash Flow per year forever = $13.6m

Assets $ 6000.800 m

Debts = $ 2000.300 m

No of Shares = 8020

1) Value of Synergy is the value of NPV of combined cash flow = Cash Flow / Rate of Return

= 13.6 m / .136 = $100m

2) Share price of Acquirer is $50 and share price of Target is $25. At $ 50 they will share equal synergy where share price of Acquirer is equals to the share price of target

in millions Acquirer Target
Assets 6000 800
Debts 2000 300
Equity (Assets - Debts) 4000 500
Shares 80 20
Price per share 50 25